# Isolated Causal Effects of Natural Language

### v1

@@ 3.3 The Fidelity-Overlap Tradeoff
$\sigma^{2}$ is a measure of the squared error of the outcome model; a smaller $\sigma^{2}$ corresponds to better model fidelity. $\nu^{2}$ is the average of the squared importance weights and so is larger in magnitude (or even unbounded) if probabilities $P(a(X)=a^{\prime}|a^{\mathsf{c}}(X))$ are more extreme. A smaller $\nu^{2}$ metric corresponds to better overlap.
With these metrics, the OVB of $\tau_{DR_{s}}$—that is, $\tau_{DR}$ using the short outcome model and short importance weight—can be bounded (Chernozhukov et al.,, 2024):
@@
where $C_{Y}$ and $C_{D}$ are user-set sensitivity parameters for the explanatory power of omitted variables toward the outcome model and importance weight.
The OVB bounds allow us to define lower and upper bounds on the isolated effect:
@@
Importantly, fidelity and overlap can be summarized by the overall robustness value of the isolated effect using the non-focal language representation $a_{s}^{\mathsf{c}}(X)$. This measure indicates how robust the effect is to OVB.
Formally, this is computed as $RV=\left|\frac{\tau_{DR_{s}}}{\sigma\nu}\right|$.
Intuitively, the robustness value can be seen as the amount of explanatory power that can be lost from approximating $a^{\mathsf{c}}(X)$ as $a_{s}^{\mathsf{c}}(X)$ before the isolated effect is no longer correct in sign (positive or negative). A larger robustness value corresponds a larger tolerance to loss of explanatory power.
We estimate $\widehat{\sigma}^{2}$, $\widehat{\nu}^{2}$, and the robustness value from the data using debiased estimators (see Appendix 1.3). We note that under this type of estimation, it is possible for the estimated $\widehat{\nu}^{2}$ to be negative; this indicates that something may have gone wrong with importance weight estimation (likely severe overlap violation).
Finally, we emphasize that while OVB may correspond to how close an effect estimate is to the ground truth, it is a complementary measure. By looking at all three metrics—$\sigma^{2}$, $\nu^{2}$, and robustness value—we can gain insight into not only which non-focal language approximations are effective but also why.
@@ 4.1 Datasets
Amazon (partially controlled setting).
The Amazon dataset (McAuley and Leskovec,, 2013) consists of reviews from the Amazon e-commerce site, each with a number of “helpful” votes. To reduce unmeasured factors in the data, we generate a new semi-synthetic outcome $Y$ by predicting the number of helpful votes as a linear function of $a(X),a_{s}^{\mathsf{c}}(X)$, then adding noise. Here, $a(X),a_{s}^{c}(X)$ encode the 10 categories from the lexicon LIWC (see Section 4.2.1) that are most predictive of vote count. These categories are binarized to take the value 1 if the category is present in the text and 0 otherwise. We note that while we control the outcome $Y$ in this data, we do not have influence over the joint distribution of $P(a(X),a_{s}^{c}(X))$.
In this partially controlled data setting, we know both (1) the true isolated effect of each of the 10 lexical categories and (2) that the outcome model $g$ can be fully learned from the text. Combined, these allow us to evaluate whether our estimator $\widehat{\tau}_{DR}$ is able to recover the true isolated effect under best-case conditions.
Semaglutide vs. Tirzepatide (SvT) (real-world setting).
Here, we consider a slightly different setting in which the intervention and outcome are both encoded in text (in contrast to the setting where the intervention is encoded in the text and the outcome is a real-valued outcome external to the text).
The SvT dataset (Dhawan et al.,, 2024) consists of posts from weight-loss communities on the social media site Reddit that mention one of two weight-loss medications: semaglutide or tirzepatide. From these posts, Dhawan et al. extracted the language-encoded binary intervention $a(X)$ (which weight-loss medication the user took) and binary outcome $Y$ (whether the user lost more than 5 percent of their starting body weight). The dataset further includes a “ground truth” causal effect from a clinical trial on the effects of semaglutide versus tirzepatide at various doses (Frías et al.,, 2021). Dhawan et al. used weight loss under 5 mg tirzepatide versus 1 mg semaglutide as the true effect. We compute confidence intervals for this true effect using information available in the clinical trial.
Using this dataset, we are able to both evaluate the validity of our isolated effect estimates in a realistic setting and see how different approximations of $a^{\mathsf{c}}(X)$ can impact our estimates.
@@ 4.2.1 Approximating non-focal language
To construct a non-focal language approximation $a_{s}^{\mathsf{c}}(X)$, we explore a number of language representations varying in complexity. In this section, we describe each representation and discuss how it might fare in the fidelity-overlap tradeoff. Implementation details can be found in Appendix 3.2.
Lexicon.
@@
One simple language representation is a vector of interpretable categories encoded by a lexicon, which maps words in its vocabulary to those categories. These categories are usually relatively few in number, so the dimensionality of the category vector is fairly low. However, lexicons are limited by their vocabulary and are also unable to capture context or sentence-level meaning. Therefore, we expect that lexicon-derived non-focal language representations may have good overlap but poor model fidelity. We use two well-known lexicons in our experiments: the human expert-designed LIWC (Pennebaker et al.,, 2015) and the semi-automatically generated Empath (Fast et al.,, 2016).
Language model embedding.
@@
Sentence embeddings extracted from transformer-based language models are among the most commonly used language representations for machine learning. These embeddings are information-rich in content and syntax, and they perform excellently on a wide variety of tasks. However, language model embeddings tend to be high-dimensional due to the size of the hidden layers in transformer architectures. As a result, embedding-derived non-focal representations may achieve good model fidelity but suffer from overlap violations. In our experiments, we use embeddings extracted from the pre-trained transformer MPNet (Song et al.,, 2020).
SenteCon.
@@
SenteCon is a language representation in which a lexicon-based layer is constructed over language model embeddings (Lin and Morency,, 2023). Like lexicon representations, a SenteCon representation consists of a vector of interpretable categories where each category is associated with a numerical weight. Because the categories are derived from an existing lexicon, the dimensionality of the SenteCon category vector should also be low. SenteCon does not rely exclusively on a pre-defined vocabulary and is able to capture sentence context. Consequently, we expect that a SenteCon-derived non-focal language representation will have reasonable model fidelity while also not being significantly affected by overlap violations. In our experiments, we use two different base lexicons for SenteCon (LIWC or Empath). We refer to these variants as SenteCon-LIWC and SenteCon-Empath.
LLM prompting.
@@
As large language model (LLM) capabilities continue to expand, it has become possible to extract attributes from a text passage simply by prompting an LLM. For instance, we might ask an LLM to tell us, based on the information contained in a paragraph, the age of the writer or if they have any health conditions. Using this form of prompting on GPT-3.5, Dhawan et al., (2024) extract a set of 10 health-related attributes from Reddit posts in the SvT dataset. Of these, we exclude 3 attributes from which weight loss can be directly computed. We treat the remaining 7 discrete variables as a type of language representation and therefore an approximation of the non-focal language.
@@ 4.2.2 Modeling and Estimation
For each non-focal language representation $a_{s}^{\mathsf{c}}(X)$, we use 5-fold cross-fitting to train an outcome model $\widehat{g}$ to predict $Y$ given $a_{s}^{\mathsf{c}}(X)$ and a classifier to predict $a(X)$ given $a_{s}^{\mathsf{c}}(X)$. We use $\widehat{P}(a(X)=a^{\prime}|a_{s}^{\mathsf{c}}(X))$ from this classifier to estimate $\widehat{\gamma}$. Within the training folds, we conduct 5-fold cross-validation to select model hyperparameters. For the Amazon dataset, we use a logistic regression classifier and a linear regression outcome model, and for the SvT dataset, we use gradient boosting models for both our classifier and outcome model. Additional model details, including libraries and hyperparameters, are available in Appendix 3.3.
With $\widehat{g}$ and $\widehat{\gamma}$ estimated, we are able to compute $\widehat{\tau}_{DR}$ on the estimation folds, which we call $\mathcal{D}_{\text{estimate}}$. When estimating the IATE, we directly use $\mathcal{D}_{\text{estimate}}$ as the source of texts $X^{*}\sim P^{*}$ for the outcome modeling term, as the target distribution is equal to the observed data distribution. When estimating the IATT, we draw $X^{*}$ from the subset of $\mathcal{D}_{\text{estimate}}$ where $a(X)=1$. We estimate the IATE for the Amazon dataset and IATT for the SvT dataset to maintain better overlap.
@@ 5 RESULTS AND DISCUSSION
(a) Isolated effect of home.
(b) Isolated effect of netspeak.
5.1 Validity of Isolated Effect Estimates
5.1.1 Amazon
@@
In the Amazon dataset, we examine the isolated effects of the 10 predictive LIWC categories used to construct the semi-synthetic outcome. We iteratively set each category to be $a(X)$ and estimate its isolated effect while using the remaining lexical categories as $a^{\mathsf{c}}(X)$. To explore the fidelity-overlap tradeoff under controlled conditions, we evaluate our isolated effect estimate and our three OVB-derived metrics—$\widehat{\sigma}^{2}$, $\widehat{\nu}^{2}$, and robustness value—under different choices of $a_{s}^{\mathsf{c}}(X)$. For each intervention, we restrict the number of remaining lexical categories used as $a_{s}^{\mathsf{c}}(X)$, beginning with 2 categories and ending with 9.
@@
Across multiple interventions (Figures 2(a), 2(b); additional results in Appendix 2), we observe that as the dimensionality (number of categories) of $a_{s}^{\mathsf{c}}(X)$ increases, so does the proximity of the isolated effect estimate to the ground truth. Moreover, the behavior of the fidelity and overlap metrics $\widehat{\sigma}^{2}$ and $\widehat{\nu}^{2}$ is also consistent with expectations. As dimensionality increases, $\widehat{\sigma}^{2}$ decreases, indicating that outcome model fidelity is improving. At the same time, $\widehat{\nu}^{2}$ increases, consistent with worsening overlap.
@@
We further see that robustness values increase with $a_{s}^{\mathsf{c}}(X)$ dimensionality, suggesting that gains in model fidelity outweigh losses in overlap. This may not be the case for all datasets, however. As this dataset does not experience significant problems with overlap (as we see from the limited range of $\widehat{\nu}^{2}$, particularly in Figure 2(a)), it seems that the outcome model performance gains are more significant in this case.
5.1.2 Semaglutide vs. Tirzepatide
In the SvT dataset, we use the intervention and outcome from Dhawan et al., (2024). We treat the Reddit post text as the non-focal language $a^{\mathsf{c}}(X)$, and we explore how each of the non-focal language representations in Section 4.2.1 impacts effect estimation.
We first observe that all of our isolated effect estimates have wide 95% confidence intervals that include both the true isolated effect and 0 (Figure 3). Looking solely at the point estimates, we see that SenteCon-Empath, SenteCon-LIWC, MPNet embeddings, and LLM prompting all yield positive isolated effect estimates—consistent with the ground truth. Of these, SenteCon-Empath comes closest to recovering the true isolated effect, but a large amount of uncertainty remains. As a result, we may not be able to use the point estimates alone to determine which representation best approximates non-focal language.
We look instead to our metrics—fidelity and overlap—where we observe interesting behavior. The high-dimensional MPNet embedding has a much larger $\widehat{\nu}^{2}$ than any other representation, and the lexicon Empath has a negative $\widehat{\nu}^{2}$; these both suggest overlap violations. Fidelity, on the other hand, is similar across all representations. In general, fidelity is expected to improve (i.e., $\sigma^{2}$ should decrease) with the dimensionality of the representation, as higher-dimensional representations are likely to contain more information; however, this may be negated by strong regularization in the outcome model.
Finally, we find that the two representations with the highest robustness values are LLM prompting and SenteCon-Empath. We note that Dhawan et al., (2024) carefully designed their LLM prompting procedure to extract discrete variables for the specific task of estimating the effect of tirzepatide versus semaglutide on weight loss. Therefore, we expect this representation to yield good results, as its robustness value reflects. Interestingly, however, the SenteCon-Empath representation—which is not specifically designed for this task—produces a similarly high robustness value, suggesting that equally effective representations can be found without requiring extensive human design effort.
@@
Figure 3: Isolated causal effect of weight-loss medication in the SvT dataset. Error bars correspond to 95% confidence intervals. The blue dotted lines surrounding the true effect mark its 95% confidence interval. Representations $a_{s}^{\mathsf{c}}(X)$ are ordered by complexity.
@@
5.2 A Closer Look at OVB and Robustness
In this section, we analyze the robustness of isolated effect estimates through the lens of OVB. When representing non-focal language, we know almost certainly that information from the text is omitted in the representation, but we do not know how much explanatory power is lost. To better understand how this loss of explanatory power affects isolated effect estimates, we focus on one representation—SenteCon-Empath in the SvT dataset—and the lower OVB bound of its associated isolated effect estimate. We recall that this bound corresponds to the lowest possible value for the effect point estimate at a given level of OVB.
In Figure 4, we plot this bound against the sensitivity parameters $C_{Y}$ and $C_{D}$, which denote the explanatory power of omitted variables toward the outcome model $g$ and the importance weight $\gamma$, respectively. This shows how the lower OVB bound changes as the hypothetical explanatory power of the omitted variables increases. We color the 0 contour red to highlight the significance of the lower OVB bound “crossing” from positive to negative as $C_{Y}$ and $C_{D}$ increase. Once the bound is negative, we can no longer be certain that the point estimate of our isolated effect is positive.
The red triangle markers on the plot provide a sense of the robustness of the isolated effect estimate. Excluding the marker labeled “masked,” each marker is an indicator of the explanatory power lost by omitting the category with that name from the SenteCon-Empath representation. [^2] These values can be computed explicitly from outcome models and importance weights estimated using both the original SenteCon-Empath representation and the version with the omitted category (see Appendix 3.4). We choose categories we believe to be relevant to the intervention and outcome. For each category omitted, we see the loss of explanatory power brings the point estimate closer to the 0 contour but is not nearly enough to cross it.
Turning then to the “masked” marker, we look at the explanatory power lost by omitting key information from the non-focal text $a^{\mathsf{c}}(X)$ itself. We create a version of the post where we mask medication type, body weights, and terms like “gain” and “loss.” This produces a much larger drop in explanatory power, but the lower OVB bound still remains positive.
From these results, we can infer that our isolated effect estimate is comparatively robust: the lower bound on the point estimate remains positive even for information loss comparable to removing relevant lexical categories or masking key information from the text.

### v3

@@ 3.3 Sensitivity to Omitted Variable Bias
where $g(a(X),a_{s}^{\mathsf{c}}(X))$ and $\gamma(a(X),a_{s}^{\mathsf{c}}(X))$ are the outcome model and importance weight that use the short non-focal language representation $a_{s}^{\mathsf{c}}(X)$. We call these the short outcome model and short importance weight. The fidelity metric indicates how close the short outcome model is to the true outcome model $g(a(X),a^{\mathsf{c}}(X))$, while the overlap metric indicates how well the overlap assumption for valid causal inference is fulfilled by the short importance weights. For both metrics, a smaller value is better.
Then the OVB of $\tau_{DR_{s}}$—that is, $\tau_{DR}$ using the short outcome model and short importance weight—is bounded:
@@
where $C_{Y}$ and $C_{D}$ are user-set sensitivity parameters for the explanatory power of omitted variables toward the outcome model and importance weight. The OVB bounds allow us to define lower and upper bounds on the isolated effect:
@@
[…]
Importantly, the fidelity-overlap tradeoff can be balanced by considering the overall robustness value of the isolated effect that uses the non-focal language representation $a_{s}^{\mathsf{c}}(X)$:
$RV=\left|\frac{\tau_{DR_{s}}}{\sigma\nu}\right|$
Intuitively, the robustness value is a measure of the effect estimate’s trustworthiness: it indicates how robust the estimate is to OVB. The robustness value can be seen as the amount of explanatory power that can be lost from approximating $a^{\mathsf{c}}(X)$ as $a_{s}^{\mathsf{c}}(X)$ before the isolated effect is no longer correct in sign (positive or negative). A larger robustness value corresponds a higher tolerance to OVB.
We estimate $\widehat{\sigma}^{2}$, $\widehat{\nu}^{2}$, and the robustness value from the data using debiased estimators (Appendix A.3). We note that under this type of estimation, it is possible for the estimated $\widehat{\nu}^{2}$ to be negative; this indicates that something may have gone wrong with importance weight estimation (potentially a severe overlap violation).
Finally, we emphasize that while OVB may correspond to how close an effect estimate is to the ground truth, it is a complementary measure. By assessing effect estimates through the lens of each metric—fidelity, overlap, and robustness value—we gain greater insight into how and why different non-focal language approximations can influence isolated effect estimation.
@@ 4.1 Datasets
Amazon (partially controlled setting). The Amazon dataset (McAuley & Leskovec, 2013) consists of reviews from the Amazon e-commerce site, each with a number of “helpful” votes. To reduce unmeasured factors in the data, we generate a new semi-synthetic outcome $Y$ by predicting the number of helpful votes as a linear function of $a(X),a_{s}^{\mathsf{c}}(X)$, then adding noise. Here, $a(X),a_{s}^{c}(X)$ encode the 10 categories from the lexicon LIWC (see Section 4.2.1) that are most predictive of vote count. These categories are binarized to take the value 1 if the category is present in the text and 0 otherwise. We note that while the semi-synthetic construction allows us to control the outcome $Y$, we do not have influence over the joint distribution $P(a(X),a_{s}^{c}(X))$.
In this partially controlled data setting, we know both (1) the true isolated effect of each of the 10 lexical categories and (2) that the outcome model $g$ can be fully learned from the text. Combined, these allow us to evaluate whether our estimator $\widehat{\tau}_{DR}$ is able to recover the true isolated effect under best-case conditions. To allow for controlled evaluation under more challenging conditions, we also generate a second semi-synthetic $Y$ where helpful votes are predicted as a non-linear function of $a(X),a_{s}^{\mathsf{c}}(X)$; this setting is discussed in Section B.2.
Semaglutide vs. Tirzepatide (SvT) (real-world setting). Here, we consider a slightly different setting in which the intervention and outcome are both encoded in text (in contrast to the setting where the intervention is encoded in the text and the outcome is a numerical value external to the text). The SvT dataset (Dhawan et al., 2024) consists of posts from weight-loss communities on the social media site Reddit that mention one of two weight-loss medications: semaglutide or tirzepatide. From these posts, Dhawan et al. extracted the language-encoded binary intervention $a(X)$ (which weight-loss medication the user took) and binary outcome $Y$ (whether the user lost more than 5 percent of their starting body weight). The dataset further includes a “ground truth” causal effect from a clinical trial on the effects of semaglutide versus tirzepatide at various doses (Frías et al., 2021). Dhawan et al. used weight loss under 5 mg tirzepatide versus 1 mg semaglutide as the true effect. We compute confidence intervals for this true effect using information available in the clinical trial.
This dataset allows us both to evaluate the validity of isolated effect estimates in a realistic setting and to assess how different approximations of $a^{\mathsf{c}}(X)$ can impact our estimates.
@@ 4.2.1 Approximating Non-Focal Language
To construct a non-focal language approximation $a_{s}^{\mathsf{c}}(X)$, we explore a number of language representations varying in complexity. In this section, we describe each representation and discuss how it might fare in the fidelity-overlap tradeoff. Implementation details can be found in Appendix C.2.
@@
Lexicon. One simple language representation is a vector of interpretable categories encoded by a lexicon, which maps words in its vocabulary to those categories. These categories are usually relatively few in number, so the dimensionality of the category vector is fairly low. However, lexicons are limited by their vocabulary and are also unable to capture context or sentence-level meaning. Therefore, we expect that lexicon-derived non-focal language representations may have good overlap but poor model fidelity. We use two well-known lexicons in our experiments: the human expert-designed LIWC (Pennebaker et al., 2015) and the semi-automatically generated Empath (Fast et al., 2016).
@@
Language model embedding. Sentence embeddings from transformer-based language models are among the most commonly used language representations for machine learning. These embeddings are information-rich in content and syntax and perform excellently on a wide variety of tasks. However, language model embeddings tend to be relatively high-dimensional compared to lexicons, and as a result, embedding-derived non-focal representations may achieve good model fidelity but suffer from overlap violations. In our experiments, we use embeddings extracted from the pre-trained transformers BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), MPNet (Song et al., 2020), and MiniLM (Wang et al., 2020). For RoBERTa and MPNet, we also use singular value decomposition (SVD) to create a lower-dimensional version of each embedding. We refer to these smaller 200-dimension representations as RoBERTa+SVD and MPNet+SVD.
@@
SenteCon. SenteCon is a language representation in which a lexicon-based layer is constructed over language model embeddings (Lin & Morency, 2023). Like lexicon representations, a SenteCon representation consists of a vector of interpretable categories where each category is associated with a numerical weight. Because the categories are derived from an existing lexicon, the dimensionality of the SenteCon category vector should also be low. SenteCon does not rely exclusively on a pre-defined vocabulary and is able to capture sentence context. Consequently, we expect that a SenteCon-derived non-focal language representation will have reasonable model fidelity while also not being significantly affected by overlap violations. In our experiments, we use two different base lexicons for SenteCon (LIWC or Empath). We refer to these variants as SenteCon-LIWC and SenteCon-Empath.
@@
LLM prompting. As large language model (LLM) capabilities continue to expand, it has become possible to extract attributes from a text passage simply by prompting an LLM. For instance, we might ask an LLM to tell us, based on the information contained in a paragraph, the age of the writer or if they have any health conditions. Using this form of prompting on GPT-3.5, Dhawan et al. (2024) extract a set of 10 health-related attributes from Reddit posts in the SvT dataset. Of these, we exclude 3 attributes from which weight loss can be directly computed. We treat the remaining 7 discrete variables as a type of language representation and therefore an approximation of the non-focal language.
@@ 4.2.2 Modeling and Estimation
For each non-focal language representation $a_{s}^{\mathsf{c}}(X)$, we use 5-fold cross-fitting to train an outcome model $\widehat{g}$ to predict $Y$ given $a_{s}^{\mathsf{c}}(X)$ and a classifier to predict $a(X)$ given $a_{s}^{\mathsf{c}}(X)$. We use $\widehat{P}(a(X)=a^{\prime}|a_{s}^{\mathsf{c}}(X))$ from this classifier to estimate $\widehat{\gamma}$. Within the training folds, we conduct 5-fold cross-validation to select model hyperparameters. For the linear-outcome case of the Amazon dataset, we use a logistic regression classifier and a linear regression outcome model (and neural networks for the nonlinear case). For the SvT dataset, we use gradient boosting models for both our classifier and outcome model. Additional model details, including libraries and hyperparameters, are available in Appendix C.3.
With $\widehat{g}$ and $\widehat{\gamma}$ estimated, we are able to compute $\widehat{\tau}_{DR}$ on the estimation folds, which we call $\mathcal{D}_{\text{estimate}}$. When estimating the IATE, we directly use $\mathcal{D}_{\text{estimate}}$ as the source of texts $X^{*}\sim P^{*}$ for the outcome modeling term, as the target distribution is equal to the observed data distribution. When estimating the IATT, we draw $X^{*}$ from the subset of $\mathcal{D}_{\text{estimate}}$ where $a(X)=1$. We estimate the IATE for the Amazon dataset and the IATT for the SvT dataset to maintain better overlap.
@@ 5 Results and Discussion
5.1 Amazon Dataset
In the Amazon dataset, we examine the isolated effects of the 10 predictive LIWC categories used to construct the semi-synthetic outcome. This section discusses the linear-function outcome, but the same trends appear in the nonlinear case (results in Section B.2).
@@ 5.1 Amazon Dataset
Across the 10 categories, we iteratively set each category to be $a(X)$ and estimate its isolated effect while using the remaining lexical categories as $a^{\mathsf{c}}(X)$. To explore the fidelity-overlap tradeoff under controlled conditions, we evaluate our isolated effect estimate and our three OVB-derived metrics—$\widehat{\sigma}^{2}$, $\widehat{\nu}^{2}$, and robustness value—under different choices of $a_{s}^{\mathsf{c}}(X)$. For each intervention, we restrict the number of remaining lexical categories used as $a_{s}^{\mathsf{c}}(X)$, beginning with 2 categories and ending with 9.
@@
Over multiple interventions (Figures 2(a), 2(b); additional results in Appendix B.1), we observe that as the dimensionality (number of categories) of $a_{s}^{\mathsf{c}}(X)$ increases, so does the proximity of the isolated effect estimate to the ground truth. Moreover, the behavior of the fidelity and overlap metrics $\widehat{\sigma}^{2}$ and $\widehat{\nu}^{2}$ is also consistent with expectations. As dimensionality increases, $\widehat{\sigma}^{2}$ decreases, indicating that outcome model fidelity is improving. At the same time, $\widehat{\nu}^{2}$ increases, consistent with worsening overlap.
@@
We further see that robustness values increase with $a_{s}^{\mathsf{c}}(X)$ dimensionality, suggesting that gains in model fidelity outweigh losses in overlap. This may not be the case for all datasets. As this dataset does not experience significant problems with overlap (as seen from the limited range of $\widehat{\nu}^{2}$, particularly in Figure 2(a)), it seems that outcome model performance gains are more significant in this case.
5.2 Semaglutide vs. Tirzepatide Dataset
@@
Figure 3: Isolated causal effect of weight-loss medication in the SvT dataset. Error bars denote 95% confidence intervals. The blue dotted lines surrounding the true effect mark its 95% confidence interval. Representations $a_{s}^{\mathsf{c}}(X)$ are ordered loosely by complexity, with less complex representations appearing closer to the top
In the SvT dataset, we use the intervention and outcome from Dhawan et al. (2024). We treat the Reddit post text as the non-focal language $a^{\mathsf{c}}(X)$, and we explore how each of the non-focal language representations in Section 4.2.1 impacts effect estimation.
We first observe that all of our isolated effect estimates have wide 95% confidence intervals that include both the true isolated effect and 0 (Figure 3). Looking solely at the point estimates, we see that almost all of the representations yield positive isolated effect estimates that are consistent with the ground truth. Of these, SenteCon-Empath comes closest to recovering the true isolated effect, with MiniLM a close second—but a large amount of uncertainty remains. As a result, we may not be able to use the point estimates alone to determine which representation best approximates non-focal language.
We look instead to fidelity and overlap, where we observe interesting behavior. The high-dimensional MPNet embedding has a much larger $\widehat{\nu}^{2}$ than any other representation, suggesting a near overlap violation. Interestingly, BERT and RoBERTa—which have the same dimensionality as MPNet—exhibit much better overlap than MPNet, possibly due to the additional optimization of MPNet for sentence-level tasks in the library used to extract its embedding. We also see that several representations, such as the lexicon Empath and the dimensionality-reduced RoBERTa+SVD, have negative $\widehat{\nu}^{2}$s. Because doubly robust estimators like the one we use for $\widehat{\nu}^{2}$ do not necessarily satisfy criteria like being non-negative in noisy settings, we hypothesize that these negative values (which are small in magnitude) may be due to noise in estimation. Fidelity, on the other hand, is much more consistent across all representations. In general, fidelity is expected to improve (i.e., $\sigma^{2}$ should decrease) with the dimensionality of the representation, as higher-dimensional representations are likely to contain more information; however, this may be negated by strong regularization in the outcome model. These results suggest that the fidelity-overlap tradeoff depends on some notion of representation complexity that may go beyond the dimensionality of the representation alone.
Finally, we find that the two representations with the highest robustness values are LLM prompting, which produces a positive but conservative effect estimate, and SenteCon-Empath, which produces an estimate very close to the true effect. MiniLM, which like SenteCon-Empath has a point estimate close to the clinical benchmark, has only a middling robustness value due to poor overlap. The robustness of LLM prompting is not unexpected: Dhawan et al. (2024) carefully designed their prompting procedure to extract discrete variables for the specific task of estimating the effect of tirzepatide versus semaglutide on weight loss, so we expect this representation to yield good results. Importantly, however, the SenteCon-Empath representation—which is not specifically designed for this task—has a similarly high robustness value, suggesting that equally effective representations can be found without requiring extensive human design effort.
Our results also illustrate the potential of dimensionality reduction methods like SVD in non-focal language approximation. Both RoBERTa and MPNet benefit from singular value thresholding across all OVB metrics: overlap improves, fidelity remains similar, and robustness value increases. Moreover, after SVD is applied, both representations’ effect point estimates move from outside the true effect confidence interval to inside the true effect confidence interval. These results suggest that dimensionality reduction can significantly improve the utility of high-dimensional non-focal language representations. Given the low computational and human overhead of these unsupervised post-processing techniques, they may often be worth trying.
@@
A closer look at OVB and robustness. While robustness values can be compared among non-focal language representations—providing some sense of how relatively robust each corresponding effect estimate is to bias—it is not immediately clear whether even the representation with the best robustness value is robust in absolute terms.
One way of understanding the scale of a robustness value is to calibrate it using the explanatory power lost when intentionally omitting variables known to have an influence on the effect estimate. We focus on one representation—SenteCon-Empath in the SvT dataset—and the lower OVB bound of its associated isolated effect estimate. We recall that this bound corresponds to the least possible value of the effect point estimate at a given level of OVB.
In Figure 4, we plot this bound against the sensitivity parameters $C_{Y}$ and $C_{D}$, which denote the explanatory power of omitted variables toward the outcome model $g$ and the importance weight $\gamma$, respectively. This contour plot shows how the lower OVB bound of the effect estimate changes as the hypothetical explanatory power of the variables omitted from the SenteCon-Empath representation increases. We color the 0 contour red to highlight the significance of the lower OVB bound “crossing” from positive to negative as $C_{Y}$ and $C_{D}$ increase. Once the bound is negative, we can no longer be certain that the point estimate of our isolated effect is positive.
We then plot four red triangles to mark the explanatory power lost by explicitly omitting the category with that name (movement, science, exercise, or healing) from the SenteCon-Empath representation. [^3] These values can be computed explicitly by re-fitting the outcome models and importance weights (Appendix C.4). We choose categories we believe to be relevant to the intervention and outcome. For each category omitted, we see the loss of explanatory power brings the point estimate closer to the 0 contour but is not nearly enough to cross it. Turning then to the masked marker, we look at the explanatory power lost by omitting key information from the non-focal text $a^{\mathsf{c}}(X)$ itself. We create a version of each Reddit post where we mask medication type, body weight, and terms like “gain” and “loss.” The resulting SenteCon-Empath representations experience a much larger drop in explanatory power, but the lower bound of the estimate still remains positive. These results suggest that the isolated effect estimate is robust (though noisy, as the wide confidence intervals may indicate), as the lower bound on the point estimate remains positive even under levels of OVB comparable to removing relevant lexical categories or masking key information from the text.